OpenStudy (jordanloveangel):

Help please anyone?!?!?1

OpenStudy (jordanloveangel):

A collection of dimes and nickels is worth $6.90. There are 96 coins in all. How many of each coin are there? dimes nickels

OpenStudy (inowalst):

Have you tried it out?

OpenStudy (alexandervonhumboldt2):

10d+5n=690 d+n=96

OpenStudy (jordanloveangel):

im confused of how 2 slove it

OpenStudy (7thgrader):

medal?

OpenStudy (alexandervonhumboldt2):

you didnt solve iit @7thGrader

OpenStudy (alexandervonhumboldt2):

no medal for you

OpenStudy (7thgrader):

or 6 dimes and 18 nickels

OpenStudy (jordanloveangel):

I NEED STEPS of how u solve it

OpenStudy (7thgrader):

well (6 x 10) + (18 x 5) = 690

OpenStudy (7thgrader):

the 10's are dimes and 5's are nickels

OpenStudy (7thgrader):

get it?

OpenStudy (7thgrader):

wait one sec i messed up

OpenStudy (7thgrader):

it's 10x + 5x

OpenStudy (7thgrader):

get it?

OpenStudy (jordanloveangel):

lil

OpenStudy (7thgrader):

(10 x 50) + (5 x 38)

OpenStudy (7thgrader):

medal?

Nnesha (nnesha):

dimes = .10 nickles = .5

OpenStudy (7thgrader):

you need 50 dimes and 38 nickels

OpenStudy (jordanloveangel):

42 dimes and 54 nickels ?

OpenStudy (7thgrader):

no 50 dimes and 38 nickles

OpenStudy (7thgrader):

what the heck where is my medal?